OpenStudy (anonymous):

please help solve the following and please show your work .5(2-x) = 1.5x+7

OpenStudy (anonymous):

multiply both sides by 10 to clear the annoying decimal. you get \[5(2-x)=15x+70\] (don't forget to multiply the 7 by 10 as well)

OpenStudy (anonymous):

then it is straighforward \[10-5x=15x+70\] \[-60=20x\] \[x=\frac{-60}{20}=-3\]